# Open-World Task and Motion Planning via Vision-Language Model Inferred Constraints

### v2

@@ A-B3 Code as Policies Implementation Details
TABLE III: Wall clock time on all tasks. We report the average wall clock time in seconds each approach took to solve each task across $10$ random seeds; the time within the 95% confidence interval is reported after the $\pm$ sign, and the average percentage of that time that was spent querying a foundation model is reported within parentheses ’()’ for every entry. We also include a row (‘Manual’) reporting results on a variant of OWL-TAMP where we manually specify ground-truth discrete and continuous constraints (and thus no need to query a foundation model for these) to serve as an approximate lower bound on the time our approach could achieve.
A-C Additional Experimental Results
In this section, we explore experimental results related to how much computation and time our various methods took on each of the simulated tasks in the RAVENS YCB-Manipulation domain from Section V. We also present a fine-grained table of soundness rate results used to construct Figure 4.
Table II shows the number of samples required by the various methods to solve each of the RAVENS-YCB Manipulation tasks. Note importantly that we do not adjust for success rates (depicted in Table I) here: many approaches that achieve a low number of samples (e.g. CaP-sample or No VLM on FruitSort) here actually do so only because they in fact fail to translate that particular task into a form that they can attempt to solve via sampling. OWL-TAMP generally does not require significantly more samples than the ‘Manual’ baseline (which serves as an approximate lower bound on the samples). In two of the tasks where it does require substantially more samples than ‘Manual’ (Mug3, SoupPour), OWL-TAMP provides discrete constraints that yield an initial plan skeleton that is incorrect, necessitating backtracking within the TAMP system, which significantly increases the number of samples.
Table III shows the wall clock time required by the various methods to solve each of the RAVENS-YCB Manipulation tasks. Note once again that we do not adjust for success rates here. In general, we see that our approach takes significantly more wall clock time than the ‘Manual’ oracle, but we also see that a substantial portion of the time taken (around 50% on average) is due to querying a foundation model. We also see that the percentage of time spent querying increases roughly with the task complexity (left to right), which is what we would expect, since in the more complex tasks with more constraints, it is more challenging to find a satisfying sample. These findings indicate that the overall runtime of our method could be significantly reduced by reducing the time taken for foundation model querying.
@@ A-C Additional Experimental Results
[…]
A-D Real Robot System Implementation
For the real-world demos, we adapt the strategy of Curtis et al. [59], which deploys TAMP without a priori object models by estimating collision, grasp, and placement affordances online. We deploy OWL-TAMP in a replanning policy [60], where at each state, the robot observes the world with its head camera, segments and estimates the object geometries, plans a course of action, and executes its plan. We used Grounding DINO [61] for object detection, segmentation, and association, RVT [62] to infer grasp affordances, TRAC-IK [63] for inverse kinematics, and cuRobo [64] for motion planning.
A-E Helper Functions for Continuous Constraint Generation.
We provide all methods (Section V) access to the following helper functions to be used towards constraining the continuous parameters used to instantiate skills. We provide the name, parameters and docstring of each function exactly as below:
@@ 85 Note that this does actually also constrain the pose’s horizontal
A-F OWL-TAMP prompting details.
As described in Section IV, our approach consists of an initial discrete constraint generation phase followed by continuous constraint generation based on a discrete partial plan skeleton with language parameters filled in.
@@
1You are an expert-level robot task planning system whose job is to help a robot accomplish the following task: ’’{task_str}’’.
2
@@
3Here is the initial predicate state (i.e., the set of all ground atoms that are true) of this task. Note that an image corresponding to the environment in this state
4is attached below:
5{initial_preds}
6
@@
7Your job is to output a sequence of ground operators (i.e., a plan) that ideally achieve the goal from this initial state.
@@
8Your plan need not be perfect, but it should capture the critical objects and actions necessary to accomplish this task (e.g.
@@
9if the task requires 4 objects being in a specific location, then you should take care to make sure the plan contains
10an action to manipulate each of the 4 objects in turn).
11
12Here are the unground operators with their descriptions.
13{nsrts_description}
14
@@
15Here are all the ground operators available to you; each operator you use in your plan must be one of these.
16{ground_operators}
17
@@
18Along with each operator in your plan, you should also output a natural language description of what that operator should
@@
19do. This description can be as detailed as you like, and should explain any details relevant to completing the particular
20ground operator successfully.
21
@@
22As an example, consider the example task ’’serve the banana inside the blue thing’’. Here, the bowl happens to be blue, and
23the initial state is:
24OnTable(banana)
25OnTable(bowl)
26And the available ground operators are:
27pick(banana)
28pick(bowl)
29pick(table)
30place_ontop(banana, bowl)
31place_inside(banana, bowl)
32place_ontop(bowl, banana)
33place_inside(bowl, banana)
34place_ontop(banana, table)
35place_inside(banana, table)
36place_ontop(bowl, table)
37place_inside(bowl, table)
38place_ontop(table, bowl)
39place_inside(table, bowl)
40
41Given this, the output should be something like:
42"""
@@
43In the initial state, there is a blue bowl on the table, and a banana atop the table. The banana is not in the bowl, and the task is to
44move the banana into the bowl.
@@
45The main actions relevant to the task are ‘pick(banana)‘ and ‘place(banana, bowl)‘. The goal involves a relationship between the banana and the bowl only.
46All other objects can be ignored.
47Plan:
@@
48pick(banana); make a stable grasp on the banana - try to make a top-down grasp for maximum likelihood of success
@@
49place(banana, bowl); place the banana stably so that it rests in the bowl - the banana is too large to fit inside the bowl if it is placed flatly: it needs to be reoriented to be upright so that it can fit into the bowl
@@
50achieve_goal(banana, bowl); the goal involves the banana being inside the bowl - this relationship is purely between the banana and bowl and doesn’t involve/require any other objects.
51"""
@@
52Notice how the plan ends in an achieve_goal operator. Every plan you output should end with such an operator, and the object arguments
@@
53to this operator (i.e., ‘(banana, bowl)‘ in this case) should be all the objects necessary to decide whether or not the goal has been achieved
@@
54(i.e., do your best not to include extraneous objects that are irrelevant to deciding whether the task goal has been achieved or not).
55
56
@@
57Please output your plan in the following format (do not include the angle brackets: those are just for illustrative purposes). Importantly, please do not list the plan with a numbered or bulleted list,
@@
58simply output each ground operator on a new line with no marking in front of the line as indicated below.
59<description of the initial state and task in your own words>
@@
60<description of which objects and actions are particularly relevant to solving the task>
@@
61<description of any challenges or other important considerations/obstacles that might arise when solving the task>
62Plan:
63<ground_operator0>; <natural language description0>
64<ground_operator1>; <natural language description1>
65...
66<ground_operatorm>; <natural language descriptionm>
@@
Note that we ask the model to output a achieve_goal operator at the end of the plan. This is used by the continuous constraint generation procedure that follows: we generate continuous constraints for this action (the natural language description of the achieve_goal operator is effectively the task’s goal description $g$) first — which effectively corresponds to generating constraints for the task goal — and then generate constraints for any previous operators in the plan by conditioning on these goal continuous constraints.
@@
1To give you an idea of what your output function should look like, here is an example function generated for the task "put the lemon on the spoon and the banana on the table", where "lemon", "spoon", "banana", and "table" are all objects in that task/scene.
2
3‘‘‘python
4def goal_check0() -> bool:
@@
5 ontop_spoon_bounds = modify_pose_bounds_to_ be_ontop_of_object(init_state, env, init_bounds, lemon.category, spoon.category)
@@
6 return position_within_bounds(lemon.pose, ontop_spoon_bounds)
7‘‘‘
8‘‘‘python
9def goal_check1() -> bool:
@@
10 ontop_table_bounds = modify_pose_bounds_to_ be_ontop_of_object(init_state, env, init_bounds, banana.category, table.category)
@@
11 return position_within_bounds(banana.pose, ontop_table_bounds)
12‘‘‘
13
@@
14Here is another example set of functions generated for the task "serve the banana inside the blue thing after drying it by placing on the plate". The initial state for this example
@@
15is shown in one of the attached images. Here, banana and bowl are both objects (the bowl happens to be blue, whereas the plate is red).
16The initial state in this example is:
17bowl: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
18banana: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
19plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
20table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
21
@@
22Importantly, notice how the goal_check function checks that the banana is ’upright’ in the bowl by checking its rotation is 90 degrees (approx. 1.57 radians)
@@
23along the roll axis. This is necessary, because the banana only fits into the bowl in this orientation, as shown in another example rendered
@@
24image attached below. Pay careful attention and think about any similar orientation constraints that might be necessary in new problems.
25
26‘‘‘python
27def goal_check0() -> bool:
@@
28 in_bowl_bounds = modify_pose_bounds_to_ be_inside_object(init_state, env, init_bounds, bowl.category)
29 banana_in_bowl_bounds = position_within_bounds(banana.pose, in_bowl_bounds)
30 is_upright = 1.4 <= abs(banana.pose.roll) <= 1.65
31 return banana_in_bowl_bounds and is_upright
32‘‘‘
@@
33Notice here that only one goal_check function was each defined, because satisfying the goal depends on all the continuous variables jointly.
@@
34Notice also that the goal_check doesn’t test for anything to do with the plate, even though "drying" the banana in the plate was important to the task. This is because - in the final state - the banana should be in the bowl (it should have previously been placed in the cup), and the ‘goal_check‘ function only operates in the final state.
35
@@
36Finally, here’s an example of constraints for a task "serve spam from its can into the cup". Here, the objects available are potted_meat_can and mug‘.
37‘‘‘python
38def goal_check0() -> bool:
@@
39 above_mug_bounds = modify_pose_bounds_to_ be_above_object(init_state, env, init_bounds, mug.category)
@@
40 above_mug = position_within_ bounds(potted_meat_can.pose, above_bowl_bounds)
41 pour_angle_sufficient = abs(potted_meat_can.pose.roll) > 1.2
42 return above_bowl and pour_angle_sufficient
43‘‘‘
@@
44Notice once again that only one goal_check function was defined.
@@
45Notice also that the function checks the roll of the potted_meat_can, because this is important to know that it has been sufficiently ’tipped-over’ such that its contents can fall from the bowl inside it into the cup.
46
@@
47Carefully consider these examples to inform your own functions for the current problem.
48...
@@
We use these same few-shot examples across all tasks without varying them. We prompt the model to output goal_check functions for the current goal and extract these. Then, for each of the previous operators in the plan with language descriptions as well as predicates that rely on the VLM for their implementation, we prompt the VLM for continuous constraints that should hold after that particular operator is executed.
A-G Code as Policies Full Prompt

### v4

@@ A.4 Additional Experimental Results
Table 3: Wall clock time on all tasks. We report the average wall clock time in seconds each approach took to solve each task across 1010 random seeds; the time within the 95% confidence interval is reported after the $\pm$ sign, and the average percentage of that time that was spent querying a foundation model is reported within parentheses ’()’ for every entry. We also include a row (‘Manual’) reporting results on a variant of OWL-TAMP where we manually specify ground-truth discrete and continuous constraints (and thus no need to query a foundation model for these) to serve as an approximate lower bound on the time our approach could achieve.
In this section, we explore experimental results related to how much computation and time our various methods took on each of the simulated tasks in the RAVENS YCB-Manipulation domain from Section 6. We also present a fine-grained table of soundness rate results used to construct Figure 4 in Table 4.
Table 2 shows the number of samples required by the various methods to solve each of the RAVENS-YCB Manipulation tasks. Note importantly that we do not adjust for success rates (depicted in Table 1) here: many approaches that achieve a low number of samples (e.g. CaP-sample or No VLM on FruitSort) here actually do so only because they in fact fail to translate that particular task into a form that they can attempt to solve via sampling. OWL-TAMP generally does not require significantly more samples than the ‘Manual’ baseline (which serves as an approximate lower bound on the samples). In two of the tasks where it does require substantially more samples than ‘Manual’ (Mug3, SoupPour), OWL-TAMP provides discrete constraints that yield an initial plan skeleton that is incorrect, necessitating backtracking within the TAMP system, which significantly increases the number of samples.
Table 3 shows the wall clock time required by the various methods to solve each of the RAVENS-YCB Manipulation tasks. Note once again that we do not adjust for success rates here. In general, we see that our approach takes significantly more wall clock time than the ‘Manual’ oracle, but we also see that a substantial portion of the time taken (around 50% on average) is due to querying a foundation model. We also see that the percentage of time spent querying increases roughly with the task complexity (left to right), which is what we would expect, since in the more complex tasks with more constraints, it is more challenging to find a satisfying sample. These findings indicate that the overall runtime of our method could be significantly reduced by reducing the time taken for foundation model querying.
@@
Table 4: Soundness rates on select methods and tasks. For tasks that require non-trivial discrete or continuous constraint generation, we present the ‘soundness rate’ (i.e., 1 - # $\frac{\text{\# false positives}}{\text{\# total trials}}$), where higher rates indicate that the method has fewer false positives.
A.5 Real Robot System Implementation
For the real-world demos, we adapted the strategy of Curtis et al. [9], which deploys TAMP without a priori object models by estimating collision, grasp, and placement affordances online. Here, we used PDDLStream and its “Adaptive” solver [14] to implement the TAMP framework and algorithm within OWL-TAMP, demonstrating OWL-TAMP’s compatibility with multiple off-the-shelf TAMP systems. We deployed OWL-TAMP in a replanning policy [15], where at each state, the robot observes the world with its head camera, segments and estimates the object geometries, plans a course of action, and executes its plan. We used Grounding DINO [39] for object detection, segmentation, and association, RVT [20] to infer grasp affordances, TRAC-IK [4] for inverse kinematics, and cuRobo [47] for motion planning.
A.6 Helper Functions for Continuous Constraint Generation.
We provide all methods (Section 6) access to the following helper functions to be used towards constraining the continuous parameters used to instantiate actions. We provide the name, parameters and docstring of each function exactly as below:
@@ 85 Note that this does actually also constrain the pose’s horizontal
A.7 OWL-TAMP prompting details.
As described in Section 5, our approach consists of an initial discrete constraint generation phase followed by continuous constraint generation based on a discrete partial plan skeleton with language parameters filled in (i.e., plan sketch).
@@
You are an expert-level robot task planning system whose job is to help a robot accomplish the following task: ’’{task_str}’’.
@@
Here is the initial predicate state (i.e., the set of all ground atoms that are true) of this task. Note that an image corresponding to the environment in this state
is attached below:
{initial_preds}
@@
Your job is to output a sequence of ground operators (i.e., a plan) that ideally achieve the goal from this initial state.
@@
Your plan need not be perfect, but it should capture the critical objects and actions necessary to accomplish this task (e.g.
@@
if the task requires 4 objects being in a specific location, then you should take care to make sure the plan contains
an action to manipulate each of the 4 objects in turn).
Here are the unground operators with their descriptions.
{nsrts_description}
@@
Here are all the ground operators available to you; each operator you use in your plan must be one of these.
{ground_operators}
@@
Along with each operator in your plan, you should also output a natural language description of what that operator should
@@
do. This description can be as detailed as you like, and should explain any details relevant to completing the particular
ground operator successfully.
@@
As an example, consider the example task ’’serve the banana inside the blue thing’’. Here, the bowl happens to be blue, and
the initial state is:
OnTable(banana)
OnTable(bowl)
And the available ground operators are:
pick(banana)
pick(bowl)
pick(table)
place_ontop(banana, bowl)
place_inside(banana, bowl)
place_ontop(bowl, banana)
place_inside(bowl, banana)
place_ontop(banana, table)
place_inside(banana, table)
place_ontop(bowl, table)
place_inside(bowl, table)
place_ontop(table, bowl)
place_inside(table, bowl)
Given this, the output should be something like:
@@
In the initial state, there is a blue bowl on the table, and a banana atop the table. The banana is not in the bowl, and the task is to
move the banana into the bowl.
@@
The main actions relevant to the task are ‘pick(banana)‘ and ‘place(banana, bowl)‘. The goal involves a relationship between the banana and the bowl only.
All other objects can be ignored.
Plan:
@@
pick(banana); make a stable grasp on the banana - try to make a top-down grasp for maximum likelihood of success
@@
place(banana, bowl); place the banana stably so that it rests in the bowl - the banana is too large to fit inside the bowl if it is placed flatly: it needs to be reoriented to be upright so that it can fit into the bowl
@@
achieve_goal(banana, bowl); the goal involves the banana being inside the bowl - this relationship is purely between the banana and bowl and doesn’t involve/require any other objects.
@@
Notice how the plan ends in an achieve_goal operator. Every plan you output should end with such an operator, and the object arguments
@@
to this operator (i.e., ‘(banana, bowl)‘ in this case) should be all the objects necessary to decide whether or not the goal has been achieved
@@
(i.e., do your best not to include extraneous objects that are irrelevant to deciding whether the task goal has been achieved or not).
@@
Please output your plan in the following format (do not include the angle brackets: those are just for illustrative purposes). Importantly, please do not list the plan with a numbered or bulleted list,
@@
simply output each ground operator on a new line with no marking in front of the line as indicated below.
<description of the initial state and task in your own words>
@@
<description of which objects and actions are particularly relevant to solving the task>
@@
<description of any challenges or other important considerations/obstacles that might arise when solving the task>
Plan:
<ground_operator0>; <natural language description0>
<ground_operator1>; <natural language description1>
<ground_operatorm>; <natural language descriptionm>
@@
Note that we ask the model to output a achieve_goal operator at the end of the plan. This is used by the continuous constraint generation procedure that follows: we generate continuous constraints for this action (the natural language description of the achieve_goal action is effectively the task’s goal description $g$) first — which effectively corresponds to generating constraints for the task goal — and then generate constraints for any previous operators in the plan by conditioning on these goal continuous constraints.
@@
To give you an idea of what your output function should look like, here is an example function generated for the task "put the lemon on the spoon and the banana on the table", where "lemon", "spoon", "banana", and "table" are all objects in that task/scene.
‘‘‘python
def goal_check0() -> bool:
@@
ontop_spoon_bounds = modify_pose_bounds_to_ be_ontop_of_object(init_state, env, init_bounds, lemon.category, spoon.category)
@@
return position_within_bounds(lemon.pose, ontop_spoon_bounds)
‘‘‘python
def goal_check1() -> bool:
@@
ontop_table_bounds = modify_pose_bounds_to_ be_ontop_of_object(init_state, env, init_bounds, banana.category, table.category)
@@
return position_within_bounds(banana.pose, ontop_table_bounds)
@@
Here is another example set of functions generated for the task "serve the banana inside the blue thing after drying it by placing on the plate". The initial state for this example
@@
is shown in one of the attached images. Here, banana and bowl are both objects (the bowl happens to be blue, whereas the plate is red).
The initial state in this example is:
bowl: Pose=RavenPose(x=-0.09269248694181442, y=-0.7042829990386963, z=0.026169249787926674, roll=0.0, pitch=-0.0, yaw=0.8605557025412023)
banana: Pose=RavenPose(x=0.17416073374449514, y=-0.33348321026557554, z=0.02017684663429707, roll=5.081222700168695e-05, pitch=0.00013538346655467005, yaw=-3.0371082921616765)
plate: Pose=RavenPose(x=-0.11636300384998322, y=-0.4429782032966614, z=0.014744692512349077, roll=7.884650441866775e-28, pitch=-7.554679105908491e-28, yaw=2.245637386214381)
table: Pose=RavenPose(x=0.0, y=-0.5, z=0.0, roll=0.0, pitch=-0.0, yaw=0.0)
@@
Importantly, notice how the goal_check function checks that the banana is upright in the bowl by checking its rotation is 90 degrees (approx. 1.57 radians)
@@
along the roll axis. This is necessary, because the banana only fits into the bowl in this orientation, as shown in another example rendered
@@
image attached below. Pay careful attention and think about any similar orientation constraints that might be necessary in new problems.
‘‘‘python
def goal_check0() -> bool:
@@
in_bowl_bounds = modify_pose_bounds_to_ be_inside_object(init_state, env, init_bounds, bowl.category)
banana_in_bowl_bounds = position_within_bounds(banana.pose, in_bowl_bounds)
is_upright = 1.4 <= abs(banana.pose.roll) <= 1.65
return banana_in_bowl_bounds and is_upright
@@
Notice here that only one goal_check function was each defined, because satisfying the goal depends on all the continuous variables jointly.
@@
Notice also that the goal_check doesnt test for anything to do with the plate, even though "drying" the banana in the plate was important to the task. This is because - in the final state - the banana should be in the bowl (it should have previously been placed in the cup), and the goal_check function only operates in the final state.
@@
Finally, heres an example of constraints for a task "serve spam from its can into the cup". Here, the objects available are potted_meat_can and mug‘.
‘‘‘python
def goal_check0() -> bool:
@@
above_mug_bounds = modify_pose_bounds_to_ be_above_object(init_state, env, init_bounds, mug.category)
@@
above_mug = position_within_ bounds(potted_meat_can.pose, above_bowl_bounds)
pour_angle_sufficient = abs(potted_meat_can.pose.roll) > 1.2
return above_bowl and pour_angle_sufficient
@@
Notice once again that only one goal_check function was defined.
@@
Notice also that the function checks the roll of the potted_meat_can, because this is important to know that it has been sufficiently tipped-over such that its contents can fall from the bowl inside it into the cup.
@@
Carefully consider these examples to inform your own functions for the current problem.
@@
We use these same few-shot examples across all tasks without varying them. We prompt the model to output goal_check functions for the current goal and extract these. Then, for each of the previous actions in the plan with language descriptions as well as predicates that rely on the VLM for their implementation, we prompt the VLM for continuous constraints that should hold after that particular action is executed.
A.8 Code as Policies Full Prompt